# Rankability-enhanced Revenue Uplift Modeling Framework for Online Marketing

### v1

@@ Abstract.
Uplift modeling has been widely employed in online marketing by predicting the response difference between the treatment and control groups, so as to identify the sensitive individuals toward interventions like coupons or discounts. Compared with traditional conversion uplift modeling, revenue uplift modeling exhibits higher potential due to its direct connection with the corporate income. However, previous works can hardly handle the continuous long-tail response distribution in revenue uplift modeling. Moreover, they have neglected to optimize the uplift ranking among different individuals, which is actually the core of uplift modeling. To address such issues, in this paper, we first utilize the zero-inflated lognormal (ZILN) loss to regress the responses and customize the corresponding modeling network, which can be adapted to different existing uplift models. Then, we study the ranking-related uplift modeling error from the theoretical perspective and propose two tighter error bounds as the additional loss terms to the conventional response regression loss. Finally, we directly model the uplift ranking error for the entire population with a listwise uplift ranking loss. The experiment results on offline public and industrial datasets validate the effectiveness of our method for revenue uplift modeling. Furthermore, we conduct large-scale experiments on a prominent online fintech marketing platform, Tencent FiT, which further demonstrates the superiority of our method in practical applications.
@@
Work done as an intern in FiT, Tencent
indicates the corresponding author, chenma@cityu.edu.hk
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Uplift Modelingccs: Applied Computing Economics
@@ 1. Introduction
Figure 1. The histogram for the revenue frequency distribution of A/B test in our deployment platform’s online fintech application scenario. Left: raw revenue values; Right: revenue values after the logarithmic transformation.
@@
The histogram of revenue frequency distribution of A/B test in Tencent’s online fintech application scenario.
@@
Though achieving acceptable performance in some idealized experiment environments or synthetic datasets, there are still some avenues to further improve the above methods. First, few of them focus on revenue uplift modeling, where the response is a continuous variable rather than a simple binary variable. This obviously increases the difficulty of precise uplift modeling. However, this type of uplift modeling problem often has higher practical value though less explored, considering that practitioners hope to increase the return (continuous response) on marketing investments in many cases. Second, most of the existing methods (Rzepakowski and Jaroszewicz, 2010; Shalit et al., 2017; Shi et al., 2019) only work well on synthetic datasets, where the range of responses is limited, which is far from the real production environments, especially in finance scenarios. In such cases, the expected response, like the mutual fund sales revenue, can range from zero/several dollars to several million or even billion dollars. If directly adopting previous uplift methods to such tasks, some outlier data points can bring huge obstacles to model learning, thus the performance can hardly be satisfactory. Besides, the distribution of revenue response can be extremely imbalanced. In the above case, the sales revenue of most users hardly exceeds 10,000 USD, while only a very tiny proportion of users may spend over 500,000 USD. These two points can be concluded as the long-tail distribution challenge, which is pretty common in online applications, like the one shown in Figure 1. Last but not least, almost all previous methods only care about the accuracy of the uplift effect prediction for different individuals, neglecting the ranking accuracy among them. That is, the individuals with higher true uplift values should obtain higher predicted uplift values compared with those lower-uplift value individuals, though the uplift value prediction is not so precise. In fact, the rankability of uplift models is a more serious and meaningful problem in many real applications, because the ultimate purpose of utilizing the uplift models is just to identify the individuals who are more susceptible to the treatment and then intervene in them differently according to the ranking of their uplift values. Moreover, this problem can be more challenging in scenarios with long-tail responses. Because in the learning phase, the model will mainly pay attention to individuals with high responses, and ignore the prediction accuracy of the individuals with lower responses as a result. Thus, the uplift ranking confusion among these latter individuals can be exacerbated.
@@
To address the aforementioned challenges, we propose a Ranking-Enhanced Revenue Uplift Modeling (RERUM) framework. First, to overcome the continuous value and long-tail distribution challenges for revenue response regression in treatment and control groups, we replace the conventional Mean-Squared-Error (MSE) loss with a ZILN loss. Meanwhile, an accompanying regression network framework is proposed. Second, we make a theoretical analysis of the uplift ranking error and propose more stringent error bounds as the additional revenue response loss terms, which can help enhance the model’s uplift ranking ability from the revenue response ranking perspective. Finally, we directly consider the uplift ranking among different individuals in the targeted population and propose a listwise ranking loss to explicitly optimize the uplift model’s rankability.
@@
  • To tackle the long-tail revenue response regression issue, we propose the ZILN loss and its corresponding model framework.
  • To mitigate the uplift modeling error, we perform a theoretical analysis and employ two tighter error bounds as response ranking loss terms to augment the above regression loss.
@@ 2. Related Works
Uplift Modeling. The uplift modeling in online marketing typically comprise of two types of problems: conversion uplift modeling and revenue uplift modeling. In conversion uplift modeling (Davis and Heller, 2017; Künzel et al., 2019; Athey and Imbens, 2015; Kane et al., 2014; Betlei et al., 2021; Knaus et al., 2021), the response label is a simple 0/1 binary variable indicating if the individual purchases the goods or takes a specific action, like clicking the advertisement. Most of previous research concentrates on this topic. Though achieving accepting performance in some scenarios, many of them (Shalit et al., 2017; Shi et al., 2019; Li and Yao, 2022; Wu et al., 2023) are dedicated to the uplift value prediction, neglecting the importance of uplift ranking among the population, which is originally the core point. Revenue uplift modeling (Gubela et al., 2020; Gubela and Lessmann, 2020; Zhou et al., 2023), however, differs from the traditional conversion uplift modeling on that its response label is a continuous variable with unlimited value range and irregular distribution, like the paying expense. This obviously introduces extra challenges for accurate uplift modeling. Unfortunately, few of works focus on the revenue uplift modeling problem, especially the rankability of revenue uplift models, though its higher application value and wider application scenarios in various online marketing platforms. In this paper, we dive into the rankability-enhanced revenue uplift modeling framework and explore its application in real-world scenarios.
@@ 3. Preliminaries
We follow the Neyman-Rubin potential outcome framework (Rubin, 2005) to formalize the revenue uplift modeling problem. In detail, we have the observed sample set $\mathcal{D}=\{(\mathbf{x}_{i},t_{i},y_{i})\}^{n}_{i=1}$. For each individual $i\in\mathcal{I}$ ($|\mathcal{I}|=n$), $y_{i}\in\mathcal{Y}\subset\mathbb{R}$ is a continuous outcome (response), $\mathbf{x}_{i}\in\mathcal{X}\subset\mathbb{R}^{d}$ is a vector of covariates, and $t_{i}\in\mathcal{T}=\{0,1\}$ denotes the treatment (with $t_{i}=0$ as the control) intervened on individual $i$. $\mathit{Y}_{i}$, $\mathit{X}_{i}$, and $\mathit{T}_{i}$ are corresponding random variables, respectively. Furthermore, $\mathit{Y}$, $\mathit{X}$, and $\mathit{T}$ are the general random variables that indicate the response, covariate, and treatment, respectively, regardless of the specific individual. Thus, the uplift effect $\tau_{i}$ of the treatment on individual $i$ is defined as:
@@
where $y_{i}^{1}$ and $y_{i}^{0}$ represent the potential responses under the corresponding treatment and control, respectively. Note that the whole population can be divided to two groups according to their received treatment: treatment group $\mathcal{D}^{t}=\{(\mathbf{x}_{i},t_{i},y_{i})|(\mathbf{x}_{i},t_{i},y_{i})\in\mathcal{D},t_{i}=1\}$, and control group $\mathcal{D}^{c}=\{(\mathbf{x}_{i},t_{i},y_{i})|(\mathbf{x}_{i},t_{i},y_{i})\in\mathcal{D},t_{i}=0\}$. $\mathbf{x}^{t/c}$ and $\mathit{X}^{t/c}$ represent the covariate vector instances of users in the treatment/control group and their corresponding random variable, respectively. The uplift modeling relies on the following four assumptions.
Assumption 1 (Stable Unit Treatment Value Assumption (SUTVA)).
This assumption requires that 1) The potential responses for any individual do not vary with the treatments assigned to other individuals, i.e., no spillover effect among the individuals; 2) For each individual, there are no different forms or versions of each treatment, which lead to different potential responses.
Assumption 2 (Ignorable Treatment Assignment (ITA)).
This assumption implies that the assignment of treatment to an individual is independent of other variables that can affect the response. It means that the treatment is either random or determined by known and measured variables, i.e., $\mathbf{\mathit{Y}}_{i}^{1},\mathbf{\mathit{Y}}_{i}^{0}\perp\!\!\!\perp\mathbf{\mathit{T}}_{i}|\mathbf{\mathit{X}}_{i}$. This is necessary to ensure that the causal effect of the treatment can be estimated without any bias. Note that in the RCT experiments, the treatment assignment is fully randomized. Thus, we can have a stronger assumption: $\mathbf{\mathit{Y}}_{i}^{1},\mathbf{\mathit{Y}}_{i}^{0}\perp\!\!\!\perp\mathbf{\mathit{T}}_{i}$.
Assumption 3 (Consistency).
This assumption states that the potential response of an individual under a specific treatment level is equal to the observed response of that individual under the same treatment level. Formally, $\forall i\in\mathcal{I},y_{i}^{\mathbf{\mathit{T}}_{i}=t_{i}}=y_{i}$.
Assumption 4 (Positivity).
Each individual $i$ has a non-zero probability of receiving any treatment, regardless of its covariates. Formally, $\forall i\in\mathcal{I},\forall\mathbf{x}_{i}\in\mathcal{X},\forall t_{i}\in\mathcal{T},0<P(\mathbf{\mathit{T}}_{i}=t_{i}|\mathbf{\mathit{X}}_{i}=\mathbf{x}_{i})<1$. This assumption ensures that the model can estimate the uplift effects of all potential treatments on all individuals, even if some treatments are rare or very unlikely to happen.
However, due to the non-simultaneous observability of potential responses, the uplift effect in the Neyman-Rubin potential outcome framework can hardly be accessed directly. An empirical alternative is the Conditional Average Treatment Effect (CATE) (Abrevaya et al., 2015) which measures the response difference between the corresponding treatment and the control, conditioned on the observed covariates. In this paper, following previous works (Betlei et al., 2021; Ma et al., 2022), we estimate the CATE from the statistical perspective as the uplift effect, that is:
@@ 4. Methodology
In this section, we first introduce the uplift estimation model framework based on the previous related models and the newly designed zero-inflated lognomal loss for the revenue response regression. Second, we theoretically analyze the uplift modeling error from the ranking perspective and derive the tighter upper error bounds as the response ranking learning objectives. Then, motivated by the listwise ranking research, we directly model the uplift ranking among different individuals in the targeted population. Finally, we conclude the overall training objective for rankability-enhanced revenue uplift modeling.
@@ 4.1. Uplift Estimation Model Framework
According to the uplift effect formulation (Eq. 2) in Sec. 3, conducting the accurate revenue response regression is undoubtedly crucial to the uplift modeling. However, previous uplift models (Shalit et al., 2017; Shi et al., 2019; Li and Yao, 2022; Wu et al., 2023) mainly rely on the MSE loss to learn response models which can hardly adapt to the continuous long-tail distribution in revenue uplift scenario, considering its sensitivity to outliers. To better achieve the goal of accurate response regression, we first study the response distribution. From Figure 1, we can have two observations: 1) small values, especially zeros occupy the main frequency of the revenue response distribution, and 2) after the logarithmic transformation, except for such small (zero) values, the rest of the data roughly follows a normal distribution.
Motivated by the above observations, we refer to the related research in customer lifetime value research (Wang et al., 2019) and propose the zero-inflated lognormal (ZILN) loss for response regression instead of the conventional MSE loss. In detail, based on the previous observation 1), the zero-inflation phenomenon, we first design a cross-entropy loss with the probability $p$ as the corresponding purchasing propensity loss to help determine if the individual will spend money. Only the individuals that do spend money (payers) will engage into the computation of the following payer expense loss. In this way, we can avoid the interference of such zero-value individuals to the positive revenue response regression. Furthermore, based on the above observation 2), we utilize the lognormal distribution with mean $\mu$ and standard deviation $\sigma$ to model the positive revenue response from such payers and correspondingly derive its negative log-likelihood as the payer expense loss for non-zero response regression. In this lognormal loss, we conduct a logarithmic transformation to reduce the skewness of the response distribution. The mathematical form of our ZILN loss is as follows:
@@
[…]
To match the realization of the above ZILN loss, we customize the uplift estimation model framework based on the previous related uplift models shown on the left side of Figure 2. As shown in the right side of Figure 2, our newly designed framework includes the representation learning module, treatment response module, control response module, etc. For a given individual, the treatment/control response module will take the corresponding representation vector from the representation learning module to predict $p^{1},\mu^{1},\sigma^{1}$ or $p^{0},\mu^{0},\sigma^{0}$, respectively, which will be further utilized to obtain regression loss via Eq. 3. Note our proposed model framework is actually agnostic to the specific uplift model selection. Any of the previously mentioned base models in Sec. 2 can be integrated with our framework. In our following implementations, we mainly adopt the deep learning-based models as the workhorse due to their strong ability to capture the complex and nonlinear relationships between covariates and responses, as well as the competitive performance proved by previous works.
@@ 4.2. Uplift Modeling Error Analysis
To restore the correct ranking among the $\tau_{i}$ for different individuals, an intuitive approach is to ensure that the predicted uplift distance $\hat{\tau}_{i}-\hat{\tau}_{j}$ between each possible individual pair $(i,j)$ is as close as possible to their true uplift distance $\tau_{i}-\tau_{j}$. Therefore, in the following part, we focus on the uplift distance prediction error $|(\hat{\tau}_{i}-\hat{\tau}_{j})-(\tau_{i}-\tau_{j})|$ and theoretically prove that conventional MSE loss for response regression is actually a loose error bound which can hardly capture the true uplift distance between individual pairs.
@@
Thus, we can obtain the discrepancy between the predicted pairwise uplift distance and the true pairwise uplift distance between individual $i$ and $j$ as follows:

### v2

@@ Abstract.
Uplift modeling has been widely employed in online marketing by predicting the response difference between the treatment and control groups, so as to identify the sensitive individuals toward interventions like coupons or discounts. Compared with traditional conversion uplift modeling, revenue uplift modeling exhibits higher potential due to its direct connection with the corporate income. However, previous works can hardly handle the continuous long-tail response distribution in revenue uplift modeling. Moreover, they have neglected to optimize the uplift ranking among different individuals, which is actually the core of uplift modeling. To address such issues, in this paper, we first utilize the zero-inflated lognormal (ZILN) loss to regress the responses and customize the corresponding modeling network, which can be adapted to different existing uplift models. Then, we study the ranking-related uplift modeling error from the theoretical perspective and propose two tighter error bounds as the additional loss terms to the conventional response regression loss. Finally, we directly model the uplift ranking error for the entire population with a listwise uplift ranking loss. The experiment results on offline public and industrial datasets validate the effectiveness of our method for revenue uplift modeling. Furthermore, we conduct large-scale experiments on a prominent online fintech marketing platform, Tencent FiT, which further demonstrates the superiority of our method in real-world applications.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671516isbn: 979-8-4007-0490-1/24/08conference: ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainccs: Information systems Uplift Modelingccs: Applied Computing Economics
@@ 1. Introduction
Figure 1. The histogram for the revenue frequency distribution of the A/B test in our deployment platform’s online fintech application scenario. Left: raw revenue values; Right: revenue values after the logarithmic transformation.
@@
The histogram for the revenue frequency distribution of the A/B test in our deployment platform’s online fintech application scenario. Left: raw revenue values; Right: revenue values after the logarithmic transformation.
@@
Though achieving acceptable performance in some idealized experiment environments or synthetic datasets, there are still some avenues to further improve the above methods. First, few of them focus on revenue uplift modeling, where the response is a continuous variable rather than a simple binary one. This obviously increases the difficulty of precise uplift modeling. Meanwhile, this type of uplift modeling problem often has higher practical value though less explored, considering that practitioners hope to increase the return (continuous response) on marketing investments in many cases. Second, most of the existing methods (Rzepakowski and Jaroszewicz, 2010; Shalit et al., 2017; Shi et al., 2019; Cui et al., 2024) only work well on synthetic datasets, where the range of responses is limited, which is far from the real production environments, especially in finance scenarios. In such cases, the expected response, like the mutual fund sales revenue, can range from zero/several dollars to several million or even billion dollars. If directly adopting previous uplift methods to such tasks, some extreme data points can bring huge obstacles to model learning, thus the performance can hardly be satisfactory. Besides, the distribution of revenue response can be extremely imbalanced. In the above case, the sales revenue of most users hardly exceeds 10,000 USD, while only a very tiny proportion of users may spend over 500,000 USD. These two points can be concluded as the long-tail distribution challenge, which is pretty common in online applications, like the one shown in Figure 1. Last but not least, almost all previous methods only care about the accuracy of the uplift effect prediction for different individuals, neglecting the ranking accuracy among them. That is, the individuals with higher true uplift values should obtain higher predicted uplift values compared with those lower-uplift value individuals, though the uplift value prediction is not so precise. In fact, the rankability of uplift models is a more serious and meaningful problem in many real-world applications, because the ultimate purpose of utilizing the uplift models is to identify the individuals who are more susceptible to the treatment and then intervene with them differently according to the ranking of their uplift values. Moreover, this problem is more challenging in scenarios with long-tail responses. Because in the learning phase, the model will mainly pay attention to individuals with high responses, and ignore the prediction accuracy of the individuals with lower responses. Thus, the uplift ranking confusion for these latter individuals can be exacerbated.
@@
To address the aforementioned challenges, we propose a Ranking-Enhanced Revenue Uplift Modeling (RERUM) framework. First, to overcome the continuous value and long-tail distribution challenges, we replace the conventional Mean-Squared-Error (MSE) loss with a ZILN loss for revenue response regression in treatment and control groups. Meanwhile, an accompanying regression network framework is proposed to adapt to ZILN loss. Second, we make a theoretical analysis of the uplift ranking error and propose more stringent error bounds as additional revenue response losses, which can help enhance the model’s uplift ranking ability from the revenue response ranking perspective. Finally, we directly consider the uplift ranking among different individuals in the targeted population and propose a listwise ranking loss to explicitly optimize the uplift model’s rankability.
@@
  • To tackle the long-tail issue in revenue response regression, we propose the ZILN loss and its corresponding model framework.
  • To mitigate the uplift modeling error, we perform a theoretical analysis and employ two tighter error bounds as response ranking losses to augment the above regression loss.
@@ 2. Related Works
Uplift Modeling. The uplift modeling in online marketing typically comprise of two types of problems: conversion uplift modeling and revenue uplift modeling. In conversion uplift modeling (Davis and Heller, 2017; Künzel et al., 2019; Kane et al., 2014; Athey and Imbens, 2015; Betlei et al., 2021), the response label is a simple 0/1 binary variable indicating if the individual purchases the goods or takes a specific action, like clicking the advertisement. Most of previous research concentrates on this topic. Though achieving accepting performance in some scenarios, many of them (Shalit et al., 2017; Shi et al., 2019; Li and Yao, 2022; Wu et al., 2023) are dedicated to the uplift value prediction, neglecting the importance of uplift ranking among the population, which is originally the core point. Revenue uplift modeling (Gubela et al., 2020; Gubela and Lessmann, 2020; Zhou et al., 2023), however, differs from the traditional conversion uplift modeling on that its response label is a continuous variable with unlimited value range and irregular distribution, like the paying expense. This obviously introduces extra challenges for accurate uplift modeling. Unfortunately, few of works focus on the revenue uplift modeling problem, especially the rankability of revenue uplift models, though its higher application value and wider application scenarios in various online marketing platforms. In this paper, we dive into the rankability-enhanced revenue uplift modeling framework and explore its application in real-world scenarios.
@@ 3. Preliminaries
We follow the Neyman-Rubin potential outcome framework (Rubin, 2005) to formulate the revenue uplift modeling problem. In detail, we have the observed sample set $\mathcal{D}=\{(\mathbf{x}_{i},t_{i},y_{i})\}^{n}_{i=1}$. For each individual $i\in\mathcal{I}$ ($|\mathcal{I}|=n$), $y_{i}\in\mathcal{Y}\subset\mathbb{R}$ is a continuous outcome (response), $\mathbf{x}_{i}\in\mathcal{X}\subset\mathbb{R}^{d}$ is a vector of covariates, and $t_{i}\in\mathcal{T}=\{0,1\}$ denotes the treatment (with $t_{i}=0$ as the control) intervened on individual $i$. $\mathit{Y}_{i}$, $\mathit{X}_{i}$, and $\mathit{T}_{i}$ are corresponding random variables, respectively. Furthermore, $\mathit{Y}$, $\mathit{X}$, and $\mathit{T}$ are the general random variables that indicate the response, covariate, and treatment, respectively, regardless of the specific individual. Thus, the uplift effect $\tau_{i}$ of the treatment on individual $i$ is defined as:
@@
where $y_{i}^{1}$ and $y_{i}^{0}$ represent the potential responses under the corresponding treatment and control, respectively. Note that the whole population can be divided to two groups according to their received treatment: treatment group $\mathcal{D}^{t}=\{(\mathbf{x}_{i},t_{i},y_{i})|(\mathbf{x}_{i},t_{i},y_{i})\in\mathcal{D},t_{i}=1\}$, and control group $\mathcal{D}^{c}=\{(\mathbf{x}_{i},t_{i},y_{i})|(\mathbf{x}_{i},t_{i},y_{i})\in\mathcal{D},t_{i}=0\}$. $\mathbf{x}^{t/c}$ and $\mathit{X}^{t/c}$ represent the covariate vector instances of users in the treatment/control group and their corresponding random variable, respectively. However, due to the non-simultaneous observability of potential responses, the uplift effect in the Neyman-Rubin potential outcome framework can hardly be accessed directly. An empirical alternative is the Conditional Average Treatment Effect (CATE) (Abrevaya et al., 2015) which measures the response difference between the corresponding treatment and the control, conditioned on the observed covariates. In this paper, following previous works (Betlei et al., 2021; Ma et al., 2022), we estimate the CATE from the statistical perspective as the uplift effect, that is:
@@ 4. Methodology
In this section, we first introduce the uplift estimation model framework based on the previous related models and the newly designed zero-inflated lognormal loss for the revenue response regression. Second, we theoretically analyze the uplift modeling error from the ranking perspective and derive the tighter upper error bounds as the response ranking learning objectives. Then, motivated by the listwise ranking research, we directly model the uplift ranking among different individuals in the targeted population. Finally, we conclude the overall training objective for rankability-enhanced revenue uplift modeling.
@@ 4.1. Uplift Estimation Model Framework
According to the uplift effect formulation (Eq. 2) in Sec. 3, conducting the accurate revenue response regression is undoubtedly crucial to the uplift modeling. However, previous uplift models (Shalit et al., 2017; Shi et al., 2019; Li and Yao, 2022; Wu et al., 2023) mainly rely on the MSE loss to learn response models which can hardly adapt to the continuous long-tail distribution in revenue uplift scenario, considering its sensitivity to outliers. To better achieve the goal of accurate response regression, we first study the response distribution. From Figure 1, we have two observations: 1) small values, especially zeros occupy the main frequency of the revenue response distribution, and 2) after the logarithmic transformation, except for such small (zero) values, the rest of the data roughly follows a normal distribution.
[…]
Motivated by the above observations, we refer to the related research in customer lifetime value research (Wang et al., 2019) and propose the zero-inflated lognormal (ZILN) loss for response regression instead of the conventional MSE loss. In detail, based on the previous observation 1), the zero-inflation phenomenon, we first design a cross-entropy loss with the probability $p$ as the corresponding purchasing propensity loss to help determine if the individual will spend money. Only the individuals that do spend money (payers) will engage into the computation of the following payer expense loss. In this way, we can avoid the interference of such zero-value individuals in the positive revenue response regression. Furthermore, based on the above observation 2), we utilize the lognormal distribution with mean $\mu$ and standard deviation $\sigma$ to model the positive revenue response from such payers and correspondingly derive its negative log-likelihood as the payer expense loss for non-zero response regression. In this lognormal loss, we conduct a logarithmic transformation to reduce the skewness of the response distribution. The mathematical form of our ZILN loss is as follows:
@@
To match the realization of the above ZILN loss, we customize the uplift estimation model framework based on the previous related uplift models shown on the left side of Figure 2. As shown in the right side of Figure 2, our newly designed framework includes the representation learning module, treatment response module, control response module, etc. For a given individual, the treatment/control response module will take the corresponding representation vector from the representation learning module to predict $p^{1},\mu^{1},\sigma^{1}$ or $p^{0},\mu^{0},\sigma^{0}$, respectively, which will be further utilized to obtain regression loss via Eq. 3. Note our proposed framework is agnostic to the specific uplift model choice; any previously mentioned base models in Sec. 2 can be integrated with our framework. In our following implementations, we mainly adopt the deep learning-based models as the workhorse due to their strong ability to capture the complex and nonlinear relations between covariates and responses, as well as the competitive performance proved by previous works.
@@ 4.2. Uplift Modeling Error Analysis
To restore the correct ranking among the $\tau_{i}$ for different individuals, an intuitive approach is to ensure that the predicted uplift distance $\hat{\tau}_{i}-\hat{\tau}_{j}$ between each possible individual pair $(i,j)$ is as close as possible to their true uplift distance $\tau_{i}-\tau_{j}$. Therefore, in the following part, we focus on the uplift distance prediction error $|(\hat{\tau}_{i}-\hat{\tau}_{j})-(\tau_{i}-\tau_{j})|$ and theoretically prove that conventional MSE loss for response regression is a loose error bound which can hardly capture the true uplift distance between individual pairs.
@@
Thus, we can obtain the discrepancy between the predicted pairwise uplift distance and the true pairwise uplift distance for individual $i$ and $j$ as follows: